### v3

# Decision Theoretic Foundations for Experiments Evaluating Human Decisions

Abstract.
How well people use information displays to make decisions is of primary interest in human-centered AI, model explainability, data visualization, and related areas. However, what constitutes a decision problem, and what is required for a study to establish that human decisions could be improved remain open to speculation. We propose a widely applicable definition of a decision problem synthesized from statistical decision theory and information economics as a standard for establishing when human decisions can be improved in HCI. We argue that to attribute loss in human performance to forms of bias, an experiment must provide participants with the information that a rational agent would need to identify the utility-maximizing decision. As a demonstration, we evaluate the extent to which recent evaluations of decision-making from the literature on AI-assisted decisions achieve this criteria. We find that only 10 (26%) of 39 studies that claim to identify biased behavior present participants with sufficient information to characterize their behavior as deviating from good decision-making in at least one treatment condition. We motivate the value of studying well-defined decision problems by describing a characterization of performance losses they allow us to conceive. In contrast, the ambiguities of a poorly communicated decision problem preclude normative interpretation. We conclude with recommendations for practice.
submissionid: 1000
1. Introduction
Providing informative interfaces to aid decision-making is a common goal in research in human-centered artificial intelligence (HCAI), visualization, and related areas. For example, visualization researchers emphasize assisting decision-making as a vital goal of data visualization (Dimara and Stasko, 2021). Similarly, in human-centered AI, a proliferation of empirical studies on human decision behavior have been deemed “necessary to evaluate the effectiveness of AI technologies in assisting decision making, but also to form a foundational understanding of how people interact with AI to make decisions” (Lai et al., 2023).
Research areas in HCI and related fields that rely heavily on decision-making lack a consensus on what a decision is, and how decisions are a distinct form of human behavior compared to tasks that are not considered decisions. Human decisions from prediction and other information displays are often evaluated in studies with the goal of identifying forms of bias: ways in which study participants’ behavior deviates from expectations of good decision-making. However, reviewing the range of tasks that get filed under decision-making, it is difficult to ascertain what common task structure makes these studies about decision-making, much less what properties a study needs to have in order for it to be capable of distinguishing human decisions as flawed.
As examples of the inconsistency and underdefinition of good decision-making in the HCI literature, colloquial definitions describe a decision task as a choice between alternatives, or a choice where the stakes are high (such as deciding which of multiple hurricane forecasts is more accurate (Padilla et al., 2015), or whether to screen child welfare cases (Zytek et al., 2021)), neither of which isolates a clear set of assumptions that must hold for a judgment to qualify as a decision. Other accounts cast decision-making as a higher-level task akin to sensemaking (Dimara and Stasko, 2021). For some, the point of studying human decisions is to compare them to a definition of idealized behavior. However, sometimes researchers define this standard in a way that is common across studies (e.g., a better forecast comes closer to the true probability of the event (Green and Chen, 2019; Fogliato et al., 2021)) while other times different definitions of the same concept are used (e.g., appropriate reliance as agreeing with an AI when it is correct and disagreeing when it’s not (Bansal et al., 2021; Buçinca et al., 2021) versus providing a higher subjective rating of trust when the model is correct (Bussone et al., 2015) versus choosing the better of the human versus the AI prediction to bet on for a future task (Dietvorst et al., 2015)). Sometimes the point is to study seemingly “subjective decisions,” like how people react to music recommendations (Kulesza et al., 2012), or decide whether a house seems to match its valuation (Prabhudesai et al., 2023), where it can seem unclear whether normative interpretations of task performance can be applied. Not only is a more coherent definition of decision needed, as recent survey papers on decision making in visualization and human-centered AI call for (Dimara and Stasko, 2021; Lai et al., 2021), but such a definition must clearly define requirements for identifying forms of bias or performance loss in human decision-making.
We argue for the identification of a minimum set of theoretical commitments for the task induced by a study of human decision-making from information displays to be well-defined. By well-defined we mean that it is possible to establish normative behavior, and consequently to conclude that observed human decisions are flawed in some way. We draw on statistical decision theory and expected utility theory (Savage, 1972; Steele and Stefánsson, 2015; Von Neumann and Morgenstern, 2007), which provide a rigorously derived and widely applicable framework for studying decision-making. Information economics formalizes information structures for decision problems(Bergemann and Morris, 2019), which may by induced by design choices about how to visualize or explain model predictions.
We contribute a widely applicable definition of a well-defined decision problem, and an argument for the value of this approach to HCAI and related research concerning decision-making from data-driven interfaces. Our first contribution is to establish and motivate a minimum set of components that must be defined for a decision problem to allow identification of an optimal decision (or set of decisions) so as to identify bias in human decisions. Using the concept of a rational Bayesian agent, we show how performance loss can only be conceived of when a study participant could in theory identify how to optimize their behavior (i.e., identify the normative decision) from the information they are provided. We find that while 39 of a sample of 46 existing studies draw conclusions about shortcomings in human decisions with prediction displays, only for 10 (26%) of these is the study task well-defined in at least one treatment condition, in the sense that participants are given sufficient information to identify, at least in principle, the best decision. We use examples to illustrate the epistemological risks to such conclusions and provide recommendations for experimenters to improve the interpretability of their study results.
2. Defining a Decision Problem
Our definition of a decision problem and corresponding standard for optimal behavior is intended for controlled evaluation of human behavior, a.k.a. normative decision research. Behavioral data (which may be produced by humans or by simulation) is collected under controlled conditions, with the goal of learning about behavior induced by the provision of information. Such studies are frequently used to describe the quality of human performance in some situation (e.g., how well people make decisions from displays in strategic settings (Zhang et al., 2023)), to rank different assistive elements by human performance (e.g, different visualizations (Correll et al., 2018) or AI explanation strategies (Liu et al., 2021)), or to test a hypothesis about how humans make decisions or what will help them do better (e.g., cognitive forcing functions will improve AI-assisted decisions (Buçinca et al., 2021)). Critically, such evaluative studies require that ground truth can be identified for any states that the study participant is queried about.
2.1. Decision Problem
Our definition of a decision problem for controlled study starts by assuming a scenario in which there is some state of the world $\theta$ which is uncertain at the time of the decision (uncertain state). The state is drawn from a state space $\Theta$ that describes the set of finite, mutually exclusive values that it can take.
The data-generating model is a distribution over values of the uncertain state and signals that provide information to the agent about the state. The data-generating model associates the state with the information that the agent is presented with (the signal) in order to inform their decision. Since the signal is correlated with the state, if the participant understands it in the context of the decision problem, they can improve their performance compared to a case where they do not have access to the signal.
When we present a person (hereafter agent) with a decision problem, we give them a choice of response (action) $a$ from some set of possible responses (action space) $A$. In decision tasks corresponding to prediction tasks, the action space is a probabilistic belief over the state space, i.e., $A=\Delta(\Theta)$. The agent faced with a decision task must consider the probability of different states induced by the signal, and choose the best action under the scoring rule, a function that assigns a score representing the quality of the action for the realized state: $S:\Theta\times A\rightarrow\mathbb{R}$.
We refer to the process that produces signals that inform on the uncertain state as a signaling policy. The relevance of the signal to the decision task is captured by the correlation between the signal and the uncertain state. More formally, a signaling policy induces an information structure that is given by a joint distribution $\pi\in\Delta(V\times\Theta)$ over signals $v\in V$ and states $\theta\in\Theta$. This joint distribution assigns to each realization $(v,\theta)\in V\times\Theta$ a probability, denoted $\pi(v,\theta)$.
If we ignore the signal $v$, $\pi$ gives rise to a probability distribution over values of the uncertain state, which we denote as the prior distribution $p\in\Delta(\Theta)$. The prior $p(\theta)$ is equivalent to $\sum\nolimits_{v\in V}\pi(v,\theta)$.
Under this definition, the formalization of the information structure $\pi$ enables a definition of normative behavior. Summarizing information about $\theta$, $\pi$ (as well as $p$) provides a basis for the experimenter to establish expectations under optimal use of the information.
2.2. Normative Behavior
Given a decision problem defined as above, we calculate the normative (“optimal”) decision by making assumptions about what it means for an agent to have coherent preferences and to optimally decide between actions under uncertainty about the outcome. Consequently, we can interpret an experiment participant’s performance as an attempt at achieving this standard, and identify sources of error (loss) in performance.
To do so, we will first assume that the agent’s preferences can be summarized by the scoring rule $S$. We assume that she will choose the action that maximizes her expected utility (score):
(1) $U(a)=S(a)$ $=\text{\bf E}_{\theta\sim q}\![{S(a,\theta)}]$
where $q$ describes the agent’s belief distribution, the probability distribution over the states that the agent believes the states of the world are drawn from, and $S(a,\theta)$ represents the score assigned by $S$.
We can define the optimal action as that which maximizes the agent’s expected utility:
(2) $a_{opt}=\operatorname*{arg\,max}_{a\in A}\text{\bf E}_{\theta\sim q}\![{S(a,\theta)}]$
2.2.1. Rational Belief Formation
To define the normative action for a decision problem, we must first characterize how we would ideally expect the agent to form the posterior beliefs $q$ about the state given the signal $v$. A theory of optimal belief construction starts by assuming that the ideal agent perfectly perceives $v$. Once perceived, how does $v$ inform the agent’s beliefs about $\theta$? Generally, we assume that the agent has some baseline set of beliefs about $\pi$ that they could use to choose an action in the absence of $v$. Encountering $v$ leads the agent to Bayesian update their beliefs.
Specifically, to calculate the optimal decision for a decision task we first define what the agent perceives the probability distribution over the state to be prior to obtaining information from $v$: $Pr(\theta)$ or $p(\theta)$ as we describe above. Whenever the signaling policy does not reveal $\pi(\theta|v)$ directly via the signal, but does inform on $\theta$, we assume that after seeing the signal, the agent updates their prior belief to a posterior belief over signals and states, using Bayes rule based on what they have learned about the data-generating model $\pi$:
(3) $q(\theta)$ $=\tfrac{Pr(v|\theta)Pr(\theta)}{Pr(v)}.$
where $Pr(v)$ is a normalizing factor. Note that the definition in Equation 3 implies that to calculate $q(\theta)$ the agent knows $\pi$:
(4) $q(\theta)=\pi(\theta|v)=\tfrac{\pi(v,\theta)}{\sum\nolimits_{\theta\in\Theta}\pi(v,\theta)}.$
Given these posterior beliefs, we use Equation 2 to identify the action that a fully rational agent would choose in order to maximize their expected utility under $S$.
2.3. Example
To make these concepts concrete, imagine a study of AI-advised decisions in which people make recidivism decisions upon seeing features of a defendant. The state is whether the person will commit another crime, i.e., recidivate, within some duration after being released, which can take the value 0 for no or 1 for yes ($\theta\in\{0,1\}$). The action space presented to the agent is whether or not to release the defendant ($a\in\{0,1\}$).
The data-generating model is the joint distribution over the state (recidivate or not) and signals comprised of defendant feature values and AI prediction information. Assume that agents are presented with a signal in the form of a set of feature values describing a defendant (age, prior arrests, prior felonies, etc.) from which they must decide on an action. Imagine that the signal also includes the prediction from a model that agents are told was trained on a historical dataset of labeled examples, along with the probability that the model’s prediction is correct. Agents can optionally use this information to inform their decision. Here, the joint distribution $\pi\in\Delta(V\times\Theta)$ assigns a probability, for each possible value of $\theta$ ({0, 1}), to each possible signal (a combination of feature values and model prediction with associated probability of correctness) used as a trial in the experiment.
If we did not show the agent any feature values at all (nor model predictions), we would expect them to make their choice based on their a priori sense of the probability that a randomly drawn defendant would recidivate. The prior distribution $p$ gives the probability that a defendant recidivates as defined by the data-generating model before viewing the signal given on an particular trial.
Experiment designs can vary in the amount of information they provide participants about $\pi$. What is critical is that a participant has some opportunity to obtain information about $\pi$. For example, imagine a version of the recidivism study in which the agent is given only the feature values and model prediction, with no information to link these to the state (i.e., no probability that the model prediction is correct, and no feedback on the realized state after their decision). Consequently, they have no opportunity to learn about $\pi$, which is necessary to establish a normative decision.
2.4. Additional Remarks on Scoring Rules
As previously mentioned, a critical yet often poorly understood component of a decision problem is a scoring rule $S$ : $A\times\Theta\to{\mathbb{R}}$ that maps the action and state to a quality or payoff. A scoring rule has two possible uses: to incentivize participants and to evaluate the quality of the decision responses.
Returning to the recidivism example, assume the researcher uses a binary scoring rule to map the ground truth realization of the state ($\theta\in{0,1}$ for not recidivate, recidivate) and the participant’s selected action from the action space ($a\in\{0,1\}$), say, releasing the defendant ($a=1$), to an outcome. For example, the agent might be scored by the following rule, which imposes the greatest penalty on false positives (i.e., acting as if the defendant will recidivate when they will not):
(5) $S(a,\theta)=\left\{\begin{array}[]{ccc}0.2&\text{if }a=0,\theta=0&\textit{release, does not recidivate }\\ -0.2&\text{if }a=0,\theta=1&\textit{release, recidivates}\\ -1&\text{if }a=1,\theta=0&\textit{not release, does not recidivate}\\ 0.5&\text{if }a=1,\theta=1&\textit{not release, recidivate}\end{array}\right.$
where the score could be in units of USD, for example.
Whether the researcher uses the same scoring rule for both incentives and evaluation versus two different rules has important implications for interpreting the results. For the recidivism example, if the researchers analyzed the collected responses using the above rule (i.e., reporting performance in payoff space) in addition to communicating it to participants for incentivization, then the utility associated with performance is defined by a single scoring rule, and the interpretation of a “good” decision is consistent across incentivization and evaluation, which is clearly beneficial for interpreting the results.
If we imagine instead that the researchers decided to consider errors on a more equal footing in analyzing their data, for example reporting the results in terms of accuracy, where each type of error is associated with the same cost, the meaning of a given action differs between incentivization and evaluation. A scoring rule induces a meaning on actions by defining a range of beliefs for which an action is optimal. Consequently, interpreting results is difficult whenever two different rules are used. One rule will by definition “misinterpret” the actions of the other by implying a different normative standard.
One way around this problem when using two different rules is to make sure that the rule used for incentivization has certain properties. Specifically, among scoring rules for beliefs, we can distinguish those that are proper scoring rules: scoring rules for which the optimal action is to predict the true distribution, i.e., $p\in\operatorname*{arg\,max}_{a\in A}S(a,p)$. When a proper scoring rule is used, the researcher can assume that an agent who understands the rule properly will report their true beliefs, as it is not possible to score higher by reporting some alternative beliefs. These beliefs can then be plugged into a different scoring rule without the risk of misinterpretation.
For any non-proper scoring rule $S:A\times\Theta\to{\mathbb{R}}$ there is an equivalent proper scoring rule $\hat{S}:\Delta(\Theta)\times\Theta\to{\mathbb{R}}$ defined by playing the optimal action under the reported belief. Formally,
(6) $\hat{S}(p,\theta)$ $=S(\operatorname*{arg\,max}\nolimits_{a\in A}S(a,p),\theta).$
Applying the construction in Equation 6 to a binary decision gives a threshold rule where for each state, all reported probabilities below a threshold receive one score and those above receive another score. Since the payoff depends only weakly on the belief, such scoring rules may be less sensitive to the agents’ understanding of the signal. For this reason, including a decision problem that directly also elicits beliefs using a proper-scoring rule is a good idea whenever eliciting responses using an action space that is not over beliefs, as it enables plugging the elicited beliefs into other scoring rules. [^1] See, e.g., Wu et al. (2023)’s analysis of (Kale et al., 2020).
The choice of possible scoring rules depends on the nature of the action space. In general, a coarser action space (i.e., with fewer choices available to participants) will produce less informative responses for the purpose of resolving observed performance relative to a normative standard. If, for example, the researcher designed the recidivism decision study to also elicit the agent’s prediction of the probability of recidivism, each trial would provide more information, and they could employ a proper scoring rule such as squared loss (i.e., Brier or quadratic rule) or log score to score the accuracy of the probabilistic beliefs. Different proper scoring rules imply different expectations about the set of possible decisions (Levinstein, 2017).
2.5. Extensions
Learning and Feedback Given a definition of normative behavior dictated by Bayesian decision theory, we would expect the decision-maker to update their beliefs given any information that informs on the joint distribution between signals and states. One source of such information is showing the agent the realization of the state after they make their choice. Imagine that participants in the recidivism decision study see feedback (i.e., are told the realized state, whether or not the defendant recidivated) after each decision they make. Assuming the signal is not revealing of the posterior probability of the state (recidivate or not), we would naturally expect the participants to learn about $\pi$ from this feedback as they complete trials, affecting their posterior beliefs (Equations 3 and 4).
Decision-Making Under Risk and Certainty The phrase “decision-making under uncertainty” is defined in the economics literature as the case where the decision-maker is not directly given the probability distribution over the state but instead must infer it. Alternatively, decision-making under risk describes the situation when this distribution is assumed to be known to the decision-maker (Steele and Stefánsson, 2015). Under such an assumption we simply replace $q$ in Equation 2 with the known distribution.
We can also easily account for so-called decision-making under certainty, where the state is deterministic [^2] For any uncertain decision, it is possible to devise an analogous set of judgments where states are deterministic and enumerate the entire state space, an approach which has been used to isolate the effect of uncertainty from the complexity of reasoning about multiple contingencies (Martínez-Marquina et al., 2019). Under certainty, we no longer need to compute expectation to calculate the optimal decision:
(7) $a_{opt}=\operatorname*{arg\,max}_{a\in A}S(a,\theta)$
3. Defining Bias in Responses to a Decision Problem
The primary motivations for using the above framework are epistemological in nature: they concern what knowledge can be gained from the results of a study. Our primary claim in this work is that to interpret the results of an experiment on human decision-making as evidence that human decisions are flawed in some way, the experiment must provide the participant with enough information to in principle identify the normative decision that will be used to judge their behavior. In other words, does the experiment give the participant enough information to align their understanding of the decision problem with its normative interpretation?
When an experiment does not provide sufficient information for the participant to solve the task in principle, there is the potential for multiplicity concerning the data-generating model. Such multiplicity has implications for interpreting the study results. As an example, consider the common approach to defining appropriate reliance on AI as agreeing with the AI when it is correct and disagreeing when it is not (Bansal et al., 2021; Gonzalez et al., 2020; Wang and Yin, 2021; Yang et al., 2020). Imagine a user receives information about the model’s overall performance unconditional on the predicted class. Consider two data-generating models that are consistent with the provided information: one where errors are always false positives, one where errors are always false negatives. Imagine a user who can only correct false positives. If we observe the user disagreeing with the model prediction some observed proportion of the time, and getting some of those instances wrong we would say they are overrelying under the first data-generating model. If they disagree with some model predictions under the second data-generating model, we would say they are underrelying. However, both are compatible with the information they were given.
3.1. Scope
It is worth making several remarks on the scope of our argument. It is important to recognize that even if the researchers do not think about the elements of their experiment in terms of the components identified above (i.e., statespace, data-generating model, signaling policy, action space, scoring rule), the design of their study still induces a decision problem whenever two simple conditions hold (Wu et al., 2023). First, there must be some uncertainty about the state (i.e., there is more than one data scenario used to generate stimuli) [^3] A single trial study where all participants view the same signal would not qualify., and second, the researchers believe it is possible to provide a scoring rule that is consistent across participants (eliminating, for example, studies of emotional responses).
It should also be noted that our argument is that the interpretation of results is restricted when a well-defined decision problem is not conveyed to participants. Researchers can certainly conduct and report on studies that do not conform to this approach, but they cannot conclude that people made bad decisions if they were not given sufficient information to know what a good decision is. For example, imagine a study comparing the decision performance of users with and without the use of an AI, where no prior is given to participants, nor a scoring rule. Imagine that the performance measured in terms of accuracy is higher access to an AI. Given an appropriately designed randomized experiment the researcher could conclude that there is a causal effect of the AI on decisions. However, they could not claim that people are making worse or biased decisions without the AI, because they would be scoring their decisions using information they never provided to participants. When we say that one condition did better, we should acknowledge that we did not give them a chance to fully understand and optimize their behavior for the problem that we analyzed.
Finally, a common misinterpretation of the use of a rational agent model is that we are implying that human behavior is expected to be perfectly rational, or approximately rational. This is not the case. While a rational agent approach formally justifies what it means to best use information to inform the selection of an action using statistical decision theory, participants in our experiment may use different strategies. The point is that without a well-defined normative response dictating what it means to best respond to the information we have provided, we are hard-pressed to evaluate participant behavior. Even if we do not expect this procedure to characterize what human agents do, it provides a basis for interpreting our results via deviations from a well-understood standard.
We we describe below, when the conditions above are met, we can characterize forms of bias as loss in decision performance according to four possible sources.
3.2. Defining Sources of Loss
Consider the process implied by definitions in Sections 2.1-2.2:
  1. (1)
    Agent obtains prior
  2. (2)
    Agent processes signal
  3. (3)
    Agent updates beliefs
  4. (4)
    Agent takes optimal action
The four steps above identify four sources of loss representing possible reasons why observed human performance in an experiment might not match the normative standard.
Prior loss (1). The agent must obtain prior beliefs that match the prior beliefs the researchers assume in analyzing the problem. Prior loss is the loss in performance due to the difference between the agent’s prior beliefs and those used by the researchers to calculate the normative standard. Explicit specification of priors is rare in decision studies used to evaluate interfaces in practice. The consequences of this vary depending on the nature of the signaling strategy. When the signal is fully revealing of the state (i.e., the posterior is a point mass indicating the state with certainty), or directly provides the posterior probability of the state, there is no need for the prior. However, whenever the signal cannot directly provide the probability of the state, the prior must be communicated for a well-defined decision problem.
Receiver loss (2). The agent processes the signal for the information it provides about the value of the uncertain state. Receiver loss is the loss due to the agent not properly extracting the information from the signal, for example, because the human visual system constrains what information is actually perceived, because the process of rendering the signal induces noise, or because participants do not understand how to read the signal.
Updating loss (3). Next the agent updates their prior beliefs with the information obtained. Updating loss is the loss due to the agent not updating their prior beliefs according to Bayes rule with the information they obtained from the signal, for example due to systematic deviations in human belief-updating relative to rational belief updating as prescribed by Bayes rule (e.g., (Camerer, 1995)).
Optimization loss (4). The agent decides by choosing a good action under their posterior beliefs. Optimization loss is the loss in performance due to not identifying the optimal action under the scoring rule. Optimization loss may include what perceptual psychologists call “lapse” responses (e.g., from fatigue or inattention), a catch-all term used to denote responses that do not necessarily reflect the underlying function capturing the person’s beliefs. Or it may include “elicitation errors,” where what participants report on the response scale does not reflect what they had in mind as a result of bias or imprecision induced by the elicitation method.
An important implication of the proper scoring rule distinction for belief reports is that in general we should not expect an agent to truthfully report their beliefs when it is not in their best interest to do so. Consequently, when non-proper rules are used for belief tasks, we cannot interpret a study participant who does not report the appropriate beliefs after examining the signal as demonstrating bias. Such “deviation” from what the researcher might expect (honest reporting) is not technically a loss, it is the expected behavior under that rule. In practice, researchers often use non-proper scoring rules when eliciting predictions; see, e.g., (Gneiting, 2011) for examples from finance where commonly used rules allow forecasters to score better by not reporting the true conditional mean of the target random variable.
Misunderstanding the decision problem. A potential contributor to all four sources of loss above is the agent misunderstanding the decision task, including the scoring rule, the statespace, and/or the action space. Even when all necessary components of a decision problem are carefully communicated, the participant might not read the instructions carefully, or might not understand them. Consequently, we should think of any tests we run to validate hypotheses about biases stemming from particular sources of performance loss as tests of the joint hypothesis comprised of the target hypothesis and a hypothesis that participants understood the task.
3.3. Examples
Researchers are limited in their ability to interpret responses to decision experiments as biased when distinct sources of performance loss that account for deviation from the normative standard cannot be conceived. We use several examples inspired by the human-centered AI and visualization literature to demonstrate why drawing conclusions about biased behavior is misleading in such scenarios.
3.3.1. AI-Assisted Recidivism Decisions
We first elaborate the recidivism decision-making study mentioned earlier. Imagine the study is designed to investigate whether people over-rely on AI assistance when deciding whether to release (i.e., grant parole to) incarcerated defendants. The researchers use a standard train-validate-split approach on a large labeled historical data set to train a model that predicts whether or not a person will recidivate. Each defendant in the dataset is represented by a set of feature values, $X=x$, which comprise part of the signal $v$ shown on each trial. Study participants are randomly assigned to either a control condition where they see the feature values along with only the model’s prediction and what is described as “the model’s confidence in its prediction, ranging between 50% and 100%,” or an AI explanation condition where they are also given additional information about how features in the data contribute to the model’s predictions, e.g., generated by LIME (Ribeiro et al., 2016). The model confidence score that is provided is the probability of a correct prediction conditional on the feature values, $Pr(correct|X=x)$.
Participants are told the AI model was trained on prior defendants whose status is known. They are told that the model has an accuracy of $80\%$ (representing the combined false positives and false negatives) on a sample of instances drawn from the test set which comprise the trials in the study.
Each participant makes parole decisions for 20 unique defendants. On each trial they are asked to choose whether or not to release the defendant ($a\in\{0,1\}$). Rather than the scoring rule given in Equation 5, imagine a simpler version that assigns different types of errors equivalent costs. Participants are given a base reward of \$$\$10USD$, but they are told they will lose \$$\$0.5USD$ from this initial endowment for every trial where they either release a person who will re-offend or continue to incarcerate a person who will not re-offend. Participants are not told how often to expect recidivism; i.e., the prior $p$ is not endowed. Participants are not given interim feedback on their decisions.
After collecting data from 500 participants, the researchers model the percentage of trials where the participants incorrectly follow the AI’s prediction. They find a statistically significant difference between the AI explanation condition and the control condition. The researchers interpret this as evidence that AI explanations promote over-reliance.
Interpretation of Results. To assess what can be said about performance loss, we first consider whether the problem is well-defined: could a rational agent could use the provided information to select the utility-maximizing decision?
The action space $A$ ($a=1$ for grant parole, $a=0$ for deny parole) is conveyed to participants, as well as the statespace $\Theta$ ($\theta=1$ for the defendant re-offending, $\theta=0$ for not). They also have knowledge of the scoring rule that assigns a score (reward amount) given the realized state and their decision.
How would a rational participant form beliefs about the posterior probability of the state using Equation 3? Participants are not provided with the prior probability of the state. This would be acceptable if the signal provided them with posterior beliefs about the uncertain state. However, the signal does not provide the required posterior probability: recall that participants receive the probability that the model is correct conditional on the features, $Pr(correct|X=x)$, but not conditional on the model’s prediction. They therefore cannot obtain sufficient information from the data-generating model $\pi$ to guide their choice of action.
Improving the experimental design. One possibility is to make the signal provide the posterior $q(\theta)$. This could be achieved if the confidence value included in the signal were to give the probability that the model is correct conditional on both the features and the prediction, e.g., $Pr(correct|X=x,y_{pred}=1)$ when the model predicts 1. In this setting, provision of the prior would not be necessary (and may confuse participants).
However, the careful reader may also have noticed another barrier to defining bias for the problem: the model’s conditional confidence is described as simply “the model’s confidence in its prediction”, making it ambiguous from the participant’s perspective whether it is an unconditional or conditional. Given that the normative response is not well-defined in the unconditional case even with the prior, this description would need specify that the confidence is conditional on the predicted label to eliminate ambiguity.
In the case where the study presented only $Pr(correct|X=x)$, not e.g., $Pr(correct|X=x,y_{pred}=1)$ when the model predicts 1, what additional information would equip the agent to optimize? If instead of the overall model accuracy over the stimuli, the experimenters provided information about the model’s error as separated rates of false positives and false negatives, the participant would be equipped with enough information to infer $Pr(v|\theta)$ in Equation 3.
It is worth noting that decision problems do not necessarily have to depend on small, discrete action spaces. Imagine another amendment to the recidivism study where instead of asking participants for a parole decision, the researchers ask them to report the probability of recidivism $Pr(\theta|v)$ (see, e.g., (Fogliato et al., 2021; Green and Chen, 2019)). Assuming the above changes are made to the design to equip participants with information to form posterior beliefs, participants could in principle report the optimal beliefs. However, if the scoring rule for a decision problem where the action space is over beliefs is not proper, we should not expect the agent to respond with their true beliefs. Optimization loss might be low, but this is not useful to the experimenters’ research questions if agents respond with a distortion of their true beliefs.
3.3.2. Election Forecast Displays
Imagine a study targeting the effect of different ways of visualizing election forecasts. The researchers develop a forecasting model to predict vote share (a percentage) by weighting and aggregating poll results from news sources for ten unique upcoming two-party state-level elections. Participants are randomly assigned to view one of three forecast displays showing the predicted vote share with uncertainty of each party’s candidate as a table, an animated graphic, or a static graphic. Assume for simplicity that all three induce informationally equivalent signals to a rational agent [^4] If any of the displays was not informationally equivalent to the others, the optimal decision may vary by display. See Wu et al. (Wu et al., 2023) for analysis of such examples using a rational agent framework..
Upon beginning the experiment, participants are randomly assigned to a display treatment and also assigned a political party preference. They are told they will view election forecasts from a model based on poll results and should imagine they live in the district where the election takes place. Each participant in the study does ten trials corresponding to the ten elections. On each trial, they are asked if they would like to vote for the candidate in the election. Voting carries no cost.
Participants are told they will win a given base reward amount plus an additional \$$\$0.25USD$ “for each trial where their candidate wins a simulated version of the election that uses the forecast model to predict how other people will vote.” The researchers generate outcomes by randomly drawing a bivariate vote share outcome from the predicted distribution under the model and generating the uncertain state $\theta$ (taking the value 0 or 1 for lose or win) by assuming the higher vote share candidate wins. Participants are not given feedback after each trial.
After collecting results for several hundred participants, the researchers fit a linear model that estimates the sensitivity of the vote decision to the closeness of the race, defined as the difference between the predicted vote share of the first candidate and 0.5. The researchers postulate that a better display induces more sensitivity. Using statistical significance testing, they reject the null hypothesis of no difference between the estimated slopes for the interaction between display type and race closeness across display types, using posthoc comparisons to declare that the animated graphic leads to better voting decisions because it induces the most sensitivity.
Interpretation of Results. To understand why the researchers’ conclusions might be misleading as a prediction of real world use of election displays, consider the decision problem from the perspective of the participants. The action space $A$ ($a=1$ for vote, $a=0$ for not vote) is conveyed to participants, and the task description also gives participants the statespace $\Theta$ ($\theta=1$ for their candidate winning, $\theta=0$ for not). Participants can infer the probability of their candidate winning from the signal showing the model predicted vote share.
What do participants know about the realized outcome of the election that the researchers describe simulating? The researchers described the use of the forecast model to simulate voting behavior, and in implemention, do not account for the participant’s vote in generating the outcome. Consequently, whether a participant votes is irrelevant to the payoff-relevant state under the scoring rule. However, the description provided to participants is ambiguous about whether the participant’s vote counts: some participants might assume, based on their knowledge of real-world elections, that it does. Hence performing well in the experiment requires being able to guess what the researchers have in mind, rather than behaving rationally in the context of the experiment.
Because, in the context of the study, voting carries no value but also no cost, we cannot conceive of the four sources of loss. There is no well-defined decision problem, and consequently, no formal justification for concluding that behavior is more biased or otherwise worse under one display condition.
Improving the experimental design. We could create a valid decision problem by changing the scoring rule so that the reward amount depends on whether one votes. For example, we could assign a cost of \$$\$0.25$ to voting:
(8) $S(a,\theta)=0.5\theta-0.25a=\left\{\begin{array}[]{ccc}0&\text{if }a=0,\theta=0&\textit{do not vote, lose election }\\ 0.5&\text{if }a=0,\theta=1&\textit{do not vote, win election}\\ -0.25&\text{if }a=1,\theta=0&\textit{vote, lose election}\\ 0.25&\text{if }a=1,\theta=1&\textit{vote, win election}\end{array}\right.$
However, the decision problem does not admit a well-defined normative response until we address the ambiguity in the description of the data-generating model. Only once participants are given a transparent description of how the realized outcome is generated from the model can we conceive of prior loss as well as updating loss, both of which should be zero if the problem is understood because the prior does not affect the judgment. We can conceive of receiver loss as the loss from participants not being able to extract the information from the signal, and optimization loss as the loss from participants not being able to choose the optimal action under their beliefs.
Of course, if we amend the description to inform participants of how the election outcome is actually generated, we should expect participants who understand the problem to never vote. Rather than informed political engagement, the proportion of participants who vote (e.g., in each treatment condition) would be better described as reflecting the level of confusion about the task. Even if the researchers did not change the description of how the realized state is generated, but designed the study to give participants feedback on the realized state after each decision, then the participants might in theory learn that their vote does not matter, which works against the researchers’ goals of presenting a situation where people could feel compelled to vote.
This example demonstrates that providing participants sufficient information to determine the utility-maximizing action is necessary but not sufficient for a meaningful decision problem in light of the research questions of interest.
3.4. Goals & Incentives in the Lab vs World
When participants are not given ample information to “solve” at least in principle the decision problem, we cannot bring well-formed expectations to the results, and should expect heterogeneity in the behaviors they exhibit. Such “heterogeneity by design” renders the results uninterpretable for evaluative purposes like identifying forms of bias.
Imagine that the researchers who designed the voting task intentionally did not include the vote in the scoring rule, and were intentionally ambiguous about the role of the vote in the data-generating model. Because the utility of voting in a real-world election is typically left implicit and likely to vary between people, the researchers might think that their results would therefore be more representative of real world voting behaviors.
Such beliefs stem from a convention among many researchers toward being optimistic about how their experimental observations relate to people’s behavior in the “real world.” Most controlled decision studies are designed to help us understand how people behave in some realistic setting that the experimental task is meant to proxy for. However, in the real world, people have goals and prior beliefs. We might not be able to perceive what utility function each individual person is using, but we can assume that behavior is goal-directed in one way or another. Savage’s axioms (Steele and Stefánsson, 2015) and the derivation of expected utility theory tell us that for behavior to be rationalizable, a person’s choices should be consistent with their beliefs about the state and the payoffs they expect under different outcomes.
Consequently, when people are in an experiment, we should not expect the analogous real world goals and beliefs for that kind of task to apply. For example, people might take actions in the real world for intrinsic value—e.g., voting in order to feel like a good citizen, or paying close attention to a defendant’s profile because the decision of whether to grant parole or not is expected to have a tremendous impact on that person’s life. It is difficult to motivate people to take actions based on intrinsic value in an experiment, unless the experiment is designed to elicit and describe those intrinsic values, in which case the normative framework here no longer applies. Hence, the idea that as experimenters we can skip fully describing the decision problem to our participants while still retaining the ability to identify their behavior as faulty or biased arises from a failure to recognize our fundamental uncertainty about how the experimental context relates to the real world.
4. Applying the Framework to HCAI Studies
We use a sample of recent evaluative studies of AI-assisted decisions to estimate how often researchers draw conclusions about flaws in human decisions in the absence of a well-defined decision problem. We sampled 46 of the studies surveyed by Lai et al. (2023) in their overview of empirical human-subjects studies on human-AI decision-making. According to their inclusion criteria, the sample contains only evaluative human-subject studies targeting a decision task in the context of classification or regression published in ACM or ACL conference venues between 2018 and 2021.
Codes. In coding the studies, we focused on three aspects of the study design and interpretation. First, we identified whether our framework is applicable: Does the task given to participants admit a ground truth state and induce uncertainty about the state, at least across participants. Next, we identified whether the authors draw conclusions about flaws or performance loss in human decisions from their results, i.e., is the study evaluating human decision-making in some way. We identified whether the study targeted a well-defined problem; i.e., whether participants in the study were provided with sufficient information for a rational agent to best respond to the task. Finally, we noted whether the researchers reported conclusions using a different scoring rule than the one provided to participants, and whether participants were incentivized using a flat scoring rule (i.e., a rule that does not distinguish between responses based on their distance to the ground truth). Note that studies that pay participants a flat reward require that researchers report conclusions using a different scoring rule. Our coding guide is in Appendix 7.2.
In some cases, coding was challenging because the paper did not provide much detail on what exactly was communicated to participants, nor provide the study interface as a supplement. We note such cases in our coding table (Appendix 7.3).
Two authors participated in the coding. After formulating the coding scheme, the authors divided the sampled studies roughly evenly, each coding independently, then discussing all challenging examples to ensure consensus. We provide the list of sampled studies with codes as Appendix.
Results. Of the 46 studies examined, 7 (15%) studied tasks to which the framework could not be applied. These included studies relegated to perceptions (e.g., of fairness, accuracy) and of subjective appraisals of music or movie recommendations or emotion. One study posed questions and drew conclusions about the diagnostic utility of AI recommendations, but elicited only perceived utility (Cai et al., 2019). We coded this example as Applicable.
All 39 of the remaining studies drew conclusions about flaws in human decisions, such as by pointing to over- or underreliance on the AI or speculating about why participants failed to get the correct answer. However, participants were only given sufficient information to identify the best decision response across all conditions in 7 of these 39 (18%), and across some conditions in another 3 of 39 (7.7%). In the latter cases, informational differences between conditions that the researchers compared could lead to different a priori expectations of decision quality, yet these differences were not accounted for in comparing participants’ behavior between conditions [^5] The calculation of separate performance benchmarks would address such ambiguity (Wu et al., 2023).. There was not enough information to determine if participants were presented with a well defined problem in 4 (10%) of the remaining studies.
Only 7 of the 39 (18%) studies to which the framework is applicable used the same scoring rule to analyze the responses as was used to incentivize participants. 25 (64%) did not, one did in at least one experiment, and 6 others (15%) did not report enough detail for us to know. Finally, 14 (36%) of the 39 used a flat rule for incentives, one did in some experiments, 16 (36%) did not, and 7 (18%) did not report enough detail for us to know.
We observed that even when authors gave participants sufficient information to identify the best response, they rarely compared the observed performance to the best attainable performance. For example, multiple papers defined a scoring rule to incentivize participants to make accurate decisions, but evaluated the responses using a simpler accuracy scale, implying that 100% accuracy was obtainable even when it would not be to a perfectly-perceiving rational agent,and potentially misinterpreting the performance of participants guided by a different scoring rule. In some cases, participants were also given sufficient information to optimize, but it did not accurately reflect the true data-generating model. The most common scenario was when the researchers gave participants information about the model’s held-out accuracy but the participants experienced instances where the model had a different accuracy. When the study design includes feedback, this leads to a contradiction between the instructions and what participants observe which makes it difficult to draw evaluative claims about behavior.
5. Recommendations for Practice
The lack of consensus around how to define a decision problem and what the minimum required components for normative analysis should be leads to research landscape in which researchers can easily be tempted to overinterpret the results of their studies, and readers may not recognize the overinterpretations. Only when researchers define the task they study as a complete decision problem can the results of experiments that study slightly different problems (e.g., use different data-generating models) be meaningfully contrasted to drive the development of theory. Adopting the framework above, which is based in established frameworks for statistical decision theory (Savage, 1972) and information design in economics (Bergemann and Morris, 2019), also stands to improve researchers’ awareness of the limitations on what be concluded from study results when a decision problem is under-communicated.
The variation induced when a decision problem is not communicated is consequential to interpretation in that it cannot be separated from variation induced by true differences in how people do a task under different conditions. Identifying the key sources of loss—much less separating them given experiment results—is a hard problem made impossible when the researcher has not clearly identified the decision problem to the study participant or the reader of a resulting research paper. Consequently, researchers wishing to design a normative decision experiment should communicate all necessary components of the decision problem both to study participants and in publications. Specifically, we recommend that researchers:
Clearly convey action and state spaces and a scoring rule. The finite action space $A$ and statespace $\Theta$ should be clearly communicated to participants, and the action space reflected in the instruments used to collect participant responses. Because the optimal action can only be determined relative to a scoring rule, participants should be given a scoring rule to reduce heterogeneity in their choice of action, even if the experimenter feels comfortable assuming that the participants will try their best and does not use the scores in assigning monetary rewards.
Provide prior and/or sufficient information about the data-generating model to calculate posterior. Sufficient information about the data-generating model must be provided to arrive at the Bayesian posterior distribution. Whenever the signal on its own does not induce a posterior, the prior should be endowed (Equation 4), and sufficient information about $\pi$ should be provided to enable estimating $Pr(v|\theta)$ and $Pr(v)$. Using presentation formats that are more likely to be automatically processed (e.g., conveying distribution information via frequency animations (Goldstein and Rothschild, 2014; hullman2015hops)) or attempting to elicit participants’ priors directly (e.g.,  (Kim et al., 2019, 2020)) may help researchers minimize prior loss.
Explicitly motivate scoring rules for incentivization and evaluation. Beyond defining the components of the target decision problem, a next step is for researchers to motivate the problem definition they use. Recent work describes the use of a rational agent framework similar to the one presented here to analyze whether a decision study is likely to provide agents with sufficient incentive to consult the signals (Wu et al., 2023). Theory around scoring rules (Li et al., 2022) and what statistical quantities are elicitable (Lambert, 2011) stand to offer prescriptive advice if incorporated in human decision research.
Compare to best attainable performance. We observed that even when researchers invest effort to design scoring rules to incentivize participants, they do not often compare the observed behavioral performance to the best attainable performance under the rule. Researchers should default to analyzing results in the space of the scoring rule used for incentives, and provide a rationale whenever they deviate. Unbiased loss estimation is a rich area for future work.
Elicit measures to estimate specific sources of loss. An important area for future work is to establish under what conditions the losses we conceptualize can be estimated. Separating losses entails adding carefully constructed questions to the experimental procedure, and/or constructing benchmarks based in rational agent performance to separate sources of loss in observed results. For example, Wu et al. (2023) use the notion of a calibrated behavioral agent along with a rational agent benchmark and baseline (expected performance of a rational agent under only the prior) to separate optimization and receiver loss.
Conduct design analysis. Researchers doing normative decision research can benefit from conducting analyses prior to running the study which simulate the idealized actions of an agent to vet the experiment design (see, e.g., Wu et al. (2023)’s rational agent-based pre-study analyses). Simulation-based design analysis forces the researcher to explicitly define the parameters of their decision problem, which can elucidate what information needs to be communicated to a participant in order to enable an optimal response. It also enables the researcher to test variations on their signaling strategy and scoring rule, which can help the researcher realize under what conditions (1) the experiment can have acceptable statistical power and (2) participants can be paid fairly. It helps the researcher choose conditions (e.g., AI explanations) that provide informative signals for the decision task, narrowing down what kinds of research questions can be answered using a particular decision problem.
6. Limitations & Future Work
Several challenges have been raised against expected utility theory as put forth by Savage (Savage, 1972); we recommend (Steele and Stefánsson, 2015) for an overview. Another critique of a normative approach dubbed “value collapse” (Nguyen, 2023) is that presenting a clear set of values as a target for behavior limits variance in ways that can be harmful. Our argument is that we should not mistake the variation that emerges from in an under-communicated decision problem as representative of behavior under a particular set of values or a particular real-world decision task. Studies aimed at understanding the range of values at play in a naturalistic decision scenario are better characterized as descriptive research. We note that while our framework is intended for normative decision research, the definitions below may also be useful if one is doing descriptive or exploratory decision research, by making clear which aspects of the decision problem are underspecified and therefore presumably the object of study. For example, a researcher might aim to characterize challenges in model-assisted decisions in a particular domain (Zytek et al., 2021), or conduct an exploratory study to elicit information about the utility function (i.e., scoring rule) that operates in a setting then communicate it back to new participants as part of the decision problem in an evaluative experiment (e.g., (Fernandes et al., 2018)).
Some have voiced pessimistic views on the value of monetary incentives for experiment participants (e.g., (Read, 2005)). These points are somewhat orthogonal to the crux of our argument, which is philosophical and concerned with what is knowable more so than the pragmatics of how much behavior changes when a problem is well-defined.
Scoring rules can sometimes be challenging for study participants to interpret; theoretical incentive compatibility does not guarantee understanding on the part of subjects. There is a rich literature discussing possible descriptions that can be used with common rules like the quadratic scoring rule (e.g., (Artinger et al., 2010)). Interfaces researchers may be in a good position to extend this literature through careful observational studies.
Ultimately, the goal of our work is to give researchers the tools to recognize that when we do experiments, we are restricted to learning about behavior in the artificial worlds we create. As much as we might want to equate our results with some real world setting used as motivation, extrapolating from the world of the controlled experiment to the real world will always be a leap of faith. It is therefore the responsibility of the experimenter to fully understand their experimental world, and, for the purposes of offering the results as a form of potentially generalizable knowledge, to give their participants a chance to do so as well.
Acknowledgements.
We thank Yifan Wu and Ziyang Guo for discussions on these topics, and Dan Weld for feedback on a draft.